OpenStudy (anonymous):

X x X -4X -13=0 HELP PLEASE

OpenStudy (anonymous):

come on i can solve this problem in my sleep

OpenStudy (anonymous):

-4x-13=0 -4x=13 divide both sides by -4..then x= -4/13

OpenStudy (anonymous):

BUT XxX ??\

OpenStudy (anonymous):

....you're talking about x times x?...it would be x^2 x^2-4x=13

OpenStudy (valpey):

Quadratic Equation: \[ax^{2}+bx+c=0\] \[x=\frac{-b\pm\sqrt{b^{2}-4ac}}{2a}\] \[x^2-4x-13=0\] \[x=\frac{4\pm\sqrt{16-52}}{2}\] \[x=\frac{4\pm\sqrt{-36}}{2}\] \[x=\frac{4\pm6i}{2}\] \[x=2\pm3i\]